1.how to find log(9.7) and antilog(-9.7)?

Grade:10

1 Answers

Ravi
9 years ago
Hello student,
You can do this question through application of derivatives: Approximation

Take f(x)= log x
f(x+ del. x) = log(x+del x)
df(x)/dx= [f(x+del.x)-f(x)]/[del.x]

So, 1/x = [log(x+del x) -log x]/del x (eqn 1)

For x=10
del x= -0.3
f(x)= log x= 1
f(x+ del. x) = log(x+del x) =log 9.7 [This is what we have to find]

Putting these in eqn 1
We get,
1/10= (log9.7 -1)/(-0.3)
log9.7 = 0.986